# Ensuring Safety in an Uncertain Environment: Constrained MDPs via Stochastic Thresholds

### v1

@@ 4.1 Reward Regret and Constraint Violation of a Pessimistic Policy
Theorem 2 (Bounds for reward regret and constraint violation of pessimistic policy).
Let $S$ be the number of states, $A$ be the number of actions, $H$ be the horizon length, $m$ be the number of constraints, $\rho$ be the Slater gap under pessimistic thresholds, $\gamma$ be the Growing-Window parameter, $T$ be the number of episodes and $\mathcal{N}$ denote the maximum number of non-zero transition probabilities across all state-action pairs. Then with probability $1-\delta$, Algorithm 2 under pessimistic thresholds obtains the following regret bounds:
@@
$\mathcal{R}_{T}(g)$ $\leq\tilde{\mathcal{O}}\left(\frac{1}{\rho}\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right),$
@@
The closest works would be Efroni et al. efroni2020exploration , who first introduce online methods that achieve $\tilde{\mathcal{O}}(\sqrt{T})$ regret and constraint violation in CMDPs; later, Ghosh et al. ghosh2024towards prove that $\tilde{\mathcal{O}}(\sqrt{T})$ regret and hard violation can be achieved by a primal-dual approach in tabular CMDPs. This paper removes the fixed-threshold assumption, and achieves comparable guarantees.
Remark 7.
Existing studies have also employed pessimistic thresholds, but the thresholds are predetermined bura2022dope . In contrast, our approach adaptively updates the thresholds during learning.
@@
We present a proof sketch here, with a detailed proof deferred to Appendix H.
@@
Decompose the sum of regret and violation scaled by dual variable, applying Lemmas 21 and 23, as follows,
$\mathcal{L}(\lambda)\leq\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})+\sum_{t=1}^{T^{\prime}}(\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}})+\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}+\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2},$ (4)
where $\mathcal{L}(\lambda)$ denotes the Lagrangian for any $0\leq\lambda\leq\rho\bf 1$ and $\eta_{\lambda}$ denotes the learning rate for dual update. The first term on the right side measures the gap between actual reward and the optimistic estimates; the second term quantifies the cumulative error between the estimated policy and the optimal policy with $\lambda_{t}$; the last two terms capture the error from updating the dual variables.
@@
Upper bound each term in the decomposition. For the first term and third term, we have $\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}\leq\frac{mH^{2}T}{2\eta_{\lambda}}$. By applying Lemma 22, we have $\left|\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})\right|\leq\tilde{\mathcal{O}}\left(\sqrt{S\mathcal{N}H^{4}T}\right)$. Bounding the second term is one of major technical contributions in this paper. We obtain the following lemma which is proved in Appendix G.
Lemma 3 (Policy optimality of pessimistic policy).
Conditioning on the success event (formally defined in Appendix D), for any episode $t\in[T^{\prime}]$, parameter $\gamma\in(0,1]$. Under pessimistic thresholds, we have
@@
Let $\{\mathcal{F}_{h}^{t}\}_{(t,h)\in[T]\times[H]}$ be the filtration generated by the state-action sequences and threshold values observed up to time step $h$ at episode $t$. Suppose that for every $(t,h)$, $\zeta_{h}^{t}(s_{h},a_{h})$ satisfies $\left|\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}[\zeta_{h}^{t}(s_{h},a_{h})]|\mathcal{F}_{h}^{t-1}\right|\leq 2H.$ Then, with probability $1-\delta$,
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}[\zeta_{h}^{t}(s_{h},a_{h})|\mathcal{F}_{h}^{t-1}]\right|\leq\tilde{\mathcal{O}}\left(\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}\right).$
@@
Combining Step 2 and Step 3, and after rearrangement, we have
$\mathcal{L}(\lambda)\lesssim$ $\left(\rho+\frac{\|\lambda\|_{2}^{2}}{\rho}\right)\sqrt{mH^{2}T}+\left(\left[(1+m\rho)+m\rho\sqrt{SA/\gamma}\right]H^{2}\sqrt{T}\right)$
$+\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA,$
Please note that this is an upper bound for the combination of reward and constraint violation. We can obtain the regret bound for reward or constraint violation through setting $\lambda=0$ or $\lambda_{i}=\rho e_{i}\mathbf{1}_{([\sum_{k=1}^{T^{\prime}}g_{i,k}]_{+}\neq 0)}$.
4.2 Reward Regret and Constraint Violations of an Optimistic Policy
When a certain degree of risk is acceptable in exchange for faster learning and higher reward potential (e.g., exploration-heavy tasks or loosely constrained resource allocation), an optimistic threshold could be adopted in Algorithm 2, by setting $\alpha_{\diamond}^{t}=\underline{\alpha}_{t}$ at every episode $t$. Similarly, we prove reward regret and constraint violations of an optimistic policy as below.
@@ 4.2 Reward Regret and Constraint Violations of an Optimistic Policy
Under the same conditions of Theorem 2, with probability $1-\delta$, Algorithm 2 under optimistic thresholds achieves the following regret bounds:
@@
This theorem shows that, even under this aggressive stance, both reward regret and constraint violation remain sublinear over episodes $T$.
@@
The proof idea is similar with the pessimistic counterpart, so we discuss the most significant differences below due to space limitation. A detailed proof is given in Appendix H.
@@
We can derive a decomposition similar to Equation (4); however, unlike the pessimistic policy case, an additional term $\sqrt{\sum_{i=1}^{m}\left(\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right)^{2}}\|\lambda\|_{2}$ appears, which measures the total estimation error for all constraints. Applying Lemma 14, we have that with probability $1-\delta$,
$\sqrt{\sum_{i=1}^{m}\left(\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right)^{2}}\leq\tilde{\mathcal{O}}\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H+1)\sqrt{m}H^{2}SA+\sqrt{mSAH^{4}T/\gamma}\right).$
@@
Optimistic policies also exhibit difference, as below.
Lemma 5 (Policy optimality of optimistic policy).
Conditioning on the success event, for any episode $t\in[T^{\prime}]$, if we adopt the optimistic thresholds, then
@@ 4.3 Safety-Performance Trade-Off
Our theoretical analysis presents a safety-performance trade-off: policies aiming for strict adherence to constraints can incur larger reward regret, while policies that aggressively maximise rewards are more prone to constraint violations. Concretely, Theorems 2 and 3 show that, compared to fixed-threshold baselines, pessimistic (strict) thresholds yield lower violations but higher reward regret, whereas optimistic thresholds exhibit the opposite behaviour.
@@
One may wonder whether it is possible to tune a hyperparameter to leverage this trade-off without resorting to two extreme cases – pessimistic or optimistic. To this end, we may introduce a hyperparameter $\xi\in[0,1]$ to interpolate between these extremes. In particular, instead of setting $\alpha_{\diamond}^{t}=\overline{\alpha}^{t}$ or $\underline{\alpha}^{t}$, we define a blended threshold
$\alpha_{\xi}^{t}=\xi\underline{\alpha}^{t}+(1-\xi)\overline{\alpha}^{t}$
at episode $t$. Varying $\xi$ adjusts how ‘strict’ or ‘relaxed’ the thresholds are, thereby controlling the balance between safety and performance. While our theorems currently focus on the two boundary cases $\xi=0$ (pessimistic) and $\xi=1$ (optimistic), it is straightforward to extend to an analogous analysis for intermediate $\xi$ values.
@@ Appendix A Summary of Notation
Constraint functions $y_{t}(s,a)=\alpha^{t}-Gq^{\pi_{t}}(p)$
@@
Occupancy measure (matrix form) $q^{\pi}(p):=[q_{h}^{\pi}(s,a;p)]_{H\times S\times A}$
Lagrangian under stochastic thresholds $\mathcal{L}(q,\lambda,\omega):=r^{T}q^{\pi}(p)+\lambda^{T}(G^{T}q^{\pi}(p)-\alpha(\omega))$
Lagrangian of Algorithm 2 $\mathcal{L^{\prime}}(q,\lambda,\omega):=\overline{r}^{T}q^{\pi}(p)+\lambda^{T}(\overline{G}^{T}q^{\pi}(p)-\alpha_{\diamond}(\omega))$
@@
Step size $\eta_{\lambda}>0$ (hyperparameter)
@@
Visitation counter $N_{t,h}(s,a)=\sum_{l=1}^{t}\mathbf{1}\{s_{h}^{l}=s,a_{h}^{l}=a\}$
@@
Optimistic estimates $\overline{r}_{t},\;\overline{g}_{t},\;\overline{\alpha}_{t}$
Pessimistic estimates $\underline{\alpha}_{t}$
@@ B.2 Lagrangian Formulation of CMDPs with Stochastic Thresholds
From these definitions, it follows that the pessimistic and optimistic scenarios define the boundaries for all feasible policies, yielding: $\mathcal{F}_{\textup{pes}}\subseteq\mathcal{F}_{q}(\omega)\subseteq\mathcal{F}_{\textup{opt}}$ with probability $1-\delta$ due to Theorem 1.
@@ Appendix C Estimators of Thresholds
For every constraint $i\in[m]$, state $s$, action $a$, step $h\in[H]$ and episode $t\in[T]$, let $(s_{h}^{l},a_{h}^{l})$ denotes the state-action pair visited in episode $l$ at step $h$; $(s_{h}^{l},a_{h}^{l},s_{h+1}^{l})$ represents the state-action pair $(s_{h}^{l},a_{h}^{l})$ is visited and the environment evolves to next state $s_{h+1}^{l}$ at step $h$ in episode $l$; $\mathbf{1}_{X}$ is the indicator function of $X$; and $N_{h}^{t}(s,a)=\sum_{l=1}^{t-1}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}$ is the total number of visits to the pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ at step $h$ before episode $t\in[T]$. First, we give the empirical averages of the thresholds as follows:
@@
where $\iota_{h}^{t}(s,a):=\sqrt{\frac{H^{2}\ln\left(\frac{2}{\delta}\right)}{2N_{t}(s,a)}}$.
@@
$\mathbb{P}\left[\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}(s,a)\right|\geq\frac{c}{N_{t}(s,a)}\right]\leq 2\exp\left(-\frac{2c^{2}}{H^{2}N_{t}(s,a)}\right).$
Setting $\delta=2\exp\left(-\frac{2c^{2}}{H^{2}N_{t}(s,a)}\right)$ and solving to find a proper value of $c$ gives the result for the threshold function. ∎
@@
where $\zeta_{h}^{t}(s,a):=\min\left\{H,\sqrt{\frac{4H^{2}\ln(mSAHT/\delta)}{\max\{1,N_{t}(s,a)\}}}\right\}$.
@@
where the inequality holds by union bound. Noticing that $\alpha_{i,h}^{t}(s,a)\leq H$, we substitute $\delta^{\prime}$ with $\delta:=\frac{\delta^{\prime}}{mSAHT}$ in $\iota_{t}(s,a)$ with an additional union bound over the possible values of $N_{t}(s,a)$. This yields $\zeta_{t}(s,a)$, and concludes the proof. ∎
@@
$\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)=\frac{\sum_{l=\max\{1,t-W_{t}+1\}}^{t}\tilde{\alpha}_{i,h}^{l}(s,a)\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}}{\max\{1,N_{h}^{(W_{t}),t+1}(s,a)\}}.$
@@
$\hat{p}_{h}^{t}(s^{\prime}\mid s,a)$ $:=\frac{\sum_{l=1}^{t}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a,s_{h+1}^{l}=s^{\prime}\}}}{\max\{1,N_{h}^{t+1}(s,a)\}}.$ (6c)

### v2

@@ 4.1 Reward Regret and Constraint Violation of a Pessimistic Policy
Let $S$, $A$, $m$, $H$, $T$ be the number of states, actions, constraints, horizon length, and episodes, respectively. Define $\rho$ as the Slater gap under pessimistic thresholds, $\gamma$ as the Growing-Window parameter, and $\mathcal{N}$ as the maximum number of non-zero transition probabilities across all state-action pairs. Then with probability at least $1-\delta$, Algorithm 2 under pessimistic thresholds achieves the following regret bounds:
@@
$\mathcal{R}_{T}(g)$ $\leq\tilde{\mathcal{O}}\left(\left(1+\frac{1}{\rho}\right)\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right),$
@@
Efroni et al. efroni2020exploration first introduce online methods that achieve $\tilde{\mathcal{O}}(\sqrt{T})$ regret and constraint violation in CMDPs; later, Ghosh et al. ghosh2024towards prove that $\tilde{\mathcal{O}}(\sqrt{T})$ regret and hard violation can be achieved by a primal-dual approach in tabular CMDPs. This paper eliminates the reliance on fixed thresholds, and achieves comparable guarantees.
@@
We present a proof sketch here, with the detailed proof provided in Appendix H.
@@
Decompose the Lagrangian $\mathcal{L}(\lambda)$, which equals the sum of the reward regret and the constraint violation scaled by dual variable as follows,
$\underbrace{\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\right)}_{(a)}+\underbrace{\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})^{\top}\lambda}_{(b)}+\underbrace{\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})}_{(c)}+\underbrace{\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}}_{(d)}+\underbrace{\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2}}_{(e)}.$ (5)
where the notation appearing in the formula is defined in Appendix G. The decomposition shows that term (a) quantifies the cumulative error between the estimated policy and the optimal policy with $\lambda_{t}$, term (b) aggregates the errors in estimating the constraints and their thresholds, term (c) measures the gap between actual reward and the optimistic estimates and terms (d)-(e) capture the error from updating the dual variables.
@@
Upper bound each term in the decomposition. Bounding term (a) is one of major technical contributions in this paper. We obtain the following lemma, with the proof deferred to Appendix G.
Lemma 3 (Policy optimality of pessimistic policies).
Conditioned on the success event (formally defined in Appendix D), for any episode $t\in[T^{\prime}]$ and parameter $\gamma\in(0,1]$. Under pessimistic thresholds, it holds
@@
Let $\{\mathcal{F}_{h}^{t}\}_{(t,h)\in[T]\times[H]}$ be the filtration generated by the state-action sequences and threshold values observed up to step $h$ at episode $t$. Suppose that for every $(t,h)$, it satisfies $\left|\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}[\zeta_{h}^{t}(s_{h},a_{h})|\mathcal{F}_{h-1}^{t}]\right|\leq 2H.$ Then, for any for any confidence parameter $\delta\in(0,1)$, with probability at least $1-\delta$, we can obtain
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\bigl{(}\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}[\zeta_{h}^{t}(s_{h},a_{h})|\mathcal{F}_{h-1}^{t}]\bigr{)}\right|\leq\tilde{\mathcal{O}}\left(\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}\right).$
In the pessimistic case, it holds that $\alpha_{i}\leq\overline{\alpha}_{t,i}$ for every constraint $i$ and episode $t$. Thus, the bound on term (b) can be obtained simply by summing the estimation errors across the constraints. For terms (c) and (d), we can derive the bounds using the definitions and Lemma 21.
@@
Combining the bounds of all terms, and after rearrangement, we get the bound for the Lagrangian $\mathcal{L}(\lambda)$. We can obtain the regret bound for reward or constraint violation through setting $\bf\lambda=0$ or $\lambda_{i}=\rho e_{i}\mathbf{1}_{\{[\sum_{t=1}^{T^{\prime}}y_{i}^{t}]_{+}>0\}}$ for each constraint $i$.
4.2 Reward Regret and Constraint Violation of an Optimistic Policy
When a certain degree of risk is acceptable in exchange for faster learning and higher reward potential (e.g., exploration-heavy tasks or loosely constrained resource allocation), an optimistic threshold can be adopted in Algorithm 2, by setting $\alpha_{\diamond}^{t}=\underline{\alpha}^{t}$ at every episode $t$. Similarly, we prove reward regret and constraint violation of an optimistic policy as below.
@@ 4.2 Reward Regret and Constraint Violation of an Optimistic Policy
Under the same conditions as Theorem 2, with probability at least $1-\delta$, Algorithm 2 under optimistic thresholds guarantees the following regret bounds:
@@
where $\rho^{\prime}$ is the Slater gap under optimistic thresholds.
This theorem shows that, even with an exploration-heavy strategy, both reward regret and constraint violation remain sublinear over $T$ episodes.
@@
The proof idea is similar to its pessimistic counterpart. We highlight the most significant differences below, given the space constraints. The detailed proof is given in Appendix H.
@@
We can derive a decomposition similar to Equation (5); however, unlike the pessimistic policy case, an additional term $\sum_{i=1}^{T^{\prime}}(\alpha^{t}-\underline{\alpha}^{t})\lambda$ arises, representing the total estimation error across all thresholds in the optimistic scenario. To derive the bound for this term, we invoke arguments analogous to those developed in Lemmas 3 and 4.
@@
Optimistic policies also exhibit distinct behaviour, as shown below.
Lemma 5 (Policy optimality of optimistic policies).
Conditioned on the success event, for any episode $t\in[T^{\prime}]$, under optimistic thresholds, the following holds:
@@ 4.3 Safety-Performance Trade-Off
Our theoretical analysis reveals a fundamental safety-performance trade-off: policies that prioritize strict constraint adherence can incur larger reward regret, while those that aggressively maximise rewards are more prone to constraint violations. Concretely, Theorems 2 and 3 demonstrate that, compared to fixed-threshold baselines, pessimistic (strict) thresholds yield lower violations but higher reward regret, whereas optimistic thresholds exhibit the opposite behaviour.
@@
One may wonder whether it is possible to tune a hyperparameter to leverage this trade-off without resorting to two extreme cases – pessimistic or optimistic. To this end, we introduce a hyperparameter $\tau\in[0,1]$ to interpolate between these extremes. In particular, instead of setting $\alpha_{\diamond}^{t}=\overline{\alpha}^{t}$ or $\underline{\alpha}^{t}$ at episode $t$, we define a blended threshold as follows:
$\alpha_{\tau}^{t}=\tau\underline{\alpha}^{t}+(1-\tau)\overline{\alpha}^{t}.$
Varying $\tau$ adjusts how ‘strict’ or ‘relaxed’ the thresholds are, thereby controlling the balance between safety and performance. While our theorems currently focus on the two boundary cases $\tau=0$ (pessimistic) and $\tau=1$ (optimistic), it is straightforward to extend to an analogous analysis for intermediate $\tau$ values.
@@ 5 Conclusion
Applicability
This paper primarily focuses on model-based approaches, specifically in the context of episodic CMDPs. While deep reinforcement learning is highly promising, it falls outside the scope of this work. Nonetheless, our work lays the key theoretical foundations for ensuring AI safety in uncertain environments.
@@ Appendix A Summary of Notation
Constraint function $y_{t}(s,a)=\alpha^{t}-Gq^{\pi_{t}}(p)$
@@
Occupancy measure (matrix form) $q^{\pi}(p)=\{q_{h}^{\pi}(s,a;p)\}_{H\times S\times A}$
Lagrangian under stochastic thresholds $\mathcal{L}(q,\lambda,\omega)=r^{T}q^{\pi}(p)+\lambda^{T}(G^{T}q^{\pi}(p)-\alpha(\omega))$
Lagrangian of Algorithm 2 $\mathcal{L^{\prime}}(q,\lambda,\omega)=\overline{r}^{T}q^{\pi}(p)+\lambda^{T}(\overline{G}^{T}q^{\pi}(p)-\alpha_{\diamond}(\omega))$
@@
Stepsize $\eta_{t}$, $\eta_{\lambda}>0$ (hyperparameter)
@@
Visitation counter $N_{h}^{t}(s,a):=\sum_{l=\max\{1,t-W_{t}+1\}}^{t}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}$
@@
Optimistic estimates $\overline{\alpha}^{t},\;\overline{r}^{t},\;\overline{g}^{t},\;\overline{p}^{t}$
Pessimistic estimates $\underline{\alpha}^{t}$
@@ B.2 Lagrangian Formulation of CMDPs with Stochastic Thresholds
From these definitions, it follows that the pessimistic and optimistic scenarios define the boundaries for all feasible policies, yielding: $\mathcal{F}_{\textup{pes}}\subseteq\mathcal{F}_{q}(\omega)\subseteq\mathcal{F}_{\textup{opt}}$ with probability at least $1-\delta$ due to Theorem 1.
@@ Appendix C Estimators of Thresholds
For every constraint $i\in[m]$, state $s$, action $a$, step $h\in[H]$ and episode $t\in[T]$, let $(s_{h}^{l},a_{h}^{l})$ denote the state-action pair visited in episode $l$ at step $h$; $(s_{h}^{l},a_{h}^{l},s_{h+1}^{l})$ represents the state-action pair $(s_{h}^{l},a_{h}^{l})$ is visited and the environment evolves to next state $s_{h+1}^{l}$ at step $h$ in episode $l$; $\mathbf{1}_{X}$ is the indicator function of $X$; and $N_{h}^{t}(s,a)=\sum_{l=1}^{t}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}$ is the total number of visits to the pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ at step $h$ before episode $t\in[T]$. First, we give the empirical averages of the thresholds as follows:
@@
where $\iota_{h}^{t}(s,a):=\sqrt{\frac{H^{2}\ln\left(\frac{2}{\delta}\right)}{2N_{h}^{t}(s,a)}}$.
@@
$\mathbb{P}\left[\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}(s,a)\right|\geq\frac{c}{N_{h}^{t}(s,a)}\right]\leq 2\exp\left(-\frac{2c^{2}}{H^{2}N_{h}^{t}(s,a)}\right).$
Setting $\delta=2\exp\left(-\frac{2c^{2}}{H^{2}N_{h}^{t}(s,a)}\right)$ and solving to find a proper value of $c$ gives the result for the threshold function. ∎
@@
where $\zeta_{h}^{t}(s,a):=\min\left\{H,\sqrt{\frac{4H^{2}\ln(mSAHT/\delta)}{\max\{1,N_{h}^{t}(s,a)\}}}\right\}$.
@@
where the inequality holds by union bound. Noticing that $\alpha_{i,h}^{t}(s,a)\leq H$, we substitute $\delta^{\prime}$ with $\delta^{\prime}=\frac{\delta}{mSAHT}$ in $\iota_{t}(s,a)$ with an additional union bound over the possible values of $N_{h}^{t}(s,a)$. This yields $\zeta_{t}(s,a)$, and concludes the proof. ∎
@@
$\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)=\frac{\sum_{l=\max\{1,t-W_{t}+1\}}^{t}\tilde{\alpha}_{i,h}^{l}(s,a)\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}}{\max\{1,N_{h}^{(W_{t}),t}(s,a)\}}.$
@@
$\hat{p}_{h}^{t}(s^{\prime}\mid s,a)$ $:=\frac{\sum_{l=1}^{t}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a,s_{h+1}^{l}=s^{\prime}\}}}{\max\{1,N_{h}^{t}(s,a)\}}.$ (7c)